OpenStudy (anonymous):

Please help me write a expression~ Write a rule so that for every input, the output is an even number!

Directrix (directrix):

Are the inputs from the set of integers? If an input were, say 1/3, then two times 1/3 = 2/3 which is not even.

OpenStudy (anonymous):

So would it be (nx2) ?

hartnn (hartnn):

yup, nx2 or 2n is absolutely correct..

OpenStudy (anonymous):

Thanks @hartnn :)

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

if you think about even function condition, f(-x)=f(x) then you will get the idea.

OpenStudy (anonymous):

hmmm that makes sense. Thanks @LeeYeongKyu !

Directrix (directrix):

So would it be (nx2) ---------- The rule would be y = 2n where n is an integer.